# Written by AAAI Press Staff1

### v2

@@ 4.3 Results on Inductivity to Partition Count
Table 6: Inductivity to unseen partition count. Here we set the target number of partitions $k=10$. I stands for inductive and T for transductive setting. In this table, the transductive version of NeuroCUT trained on $k=10$ serves as a point of reference when assessing the quality of the inductive version of NeuroCUT.
@@
[…]
Figure 2: Results on the initial warm start and the final cut values obtained by NeuroCUT at $k{=}10$. It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function. Subsequently, there is a significant difference between the initial and final cut values.
As detailed in Section 3.2, the decoupling of parameter size and the number of partitions allows NeuroCUT to generalize effectively to an unseen number of partitions. In this section, we empirically analyze the generalization performance of NeuroCUT to an unknown number of partitions. We compare it with the non-neural baseline K-means, as the neural methods like GAP, DMon, Ortho, and MinCutPool cannot be employed to infer on an unseen number of partitions. In Table 6, we present the results on inductivity. Specifically, we trained NeuroCUT on different partition sizes ($k=5$ and $k=8$) and tested it on an unseen partition size $k=10$. The inductive version, referred to as NeuroCUT-I [^2] For clarity purposes, in this experiment we use NeuroCUT-T to refer to the transductive setting whose results are presented in Sec.  4.2, obtains high-quality results on different datasets. Note that while the non-neural method such as K-means needs to be re-run for the unseen partitions, NeuroCUT only needs to perform forward pass to produce the results on an unseen partition size.
4.4 Ablation Studies
Figure 3: Node Selection in Phase 1 with Heuristic vs Random: Relative % improvement (gain) in cut values obtained by NeuroCUT when using node selection heuristic against random node selection for $k=5$. In most cases, our heuristic finds significantly better cuts than a random node selection procedure.
[…]
5 Conclusion
In this work, we study the problem of graph partitioning with node features. Existing neural methods for addressing this problem require the target objective to be differentiable and necessitate prior knowledge of the number of partitions. In this paper, we introduced NeuroCUT, a framework to effectively address the graph partitioning problem with node features. NeuroCUT tackles these challenges using a reinforcement learning-based approach that can adapt to any target objective function. Further, attributed to its decoupled parameter space and partition count, NeuroCUT can generalize to an unseen number of partitions. The efficacy of our approach is empirically validated through an extensive evaluation on four datasets, four graph partitioning objectives and diverse partition counts. Notably, our method shows significant performance gains when compared to the state-of-the-art techniques, proving its competence in both inductive and transductive settings. In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
@@ Appendix A Appendix
A.1 Additional Related Work
In recent years there has been a growing interest in using machine learning techniques to design heuristics for NP-Hard combinatorial problems on graphs. This line of work is motivated from the observation that traditional algorithmic approaches are limited to using the same heuristic irrespective of the underlying data distribution. Neural approaches, instead, focus on learning the heuristic as a function of the data without deriving explicit algorithms (duong2023deep; jung2022learning; nazi2019gap). Significant advances have been made in learning to solve a variety of combinatorial optimization problems such as Set Cover, Travelling Salesman Problem, Influence Maximization etc. (khalil2017learning; gcomb; kool2018attention; joshi2019efficient). The neural architectural designs in these methods are generally suitable for a specific set of CO problems (mazyavkina2021reinforcement) of similar nature. For instance, AM  (kool2018attention) and NeuroLKH (xin2021neurolkh) are targeted towards routing problems. GCOMB (gcomb) and LeNSE (ireland2022lense) aim to solve problems that are submodular in nature. Owing to these assumptions, the architectural designs in these works are not suitable for our problem of $k$-way inductive graph partitioning with versatile partitioning objectives.
A.2 Clustering for Initialization
@@
As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions. Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its raw node features and positional representation i.e $\mathbf{emb_{init}}(u)\;\forall u\in\mathcal{V}$ based upon eq. 8. Further, we used $L_{\inf}$ norm as the distance metric for clustering.
A.3 Dataset Statistics
Table  7 describes the statistics of datasets used in our work. For all datasets we use their largest connected component.
Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Number of features
Cora 2485 5069 1433
CiteSeer 2120 3679 3703
Harbin 6598 10492 2
Actor 6198 14879 931
Table 7: Dataset Statistics
A.4 Time complexity of NeuroCUT
We derive the time complexity of forward pass of NeuroCUT as discussed in sec. 3.4 in main paper.
@@ A.4 Time complexity of NeuroCUT
(1) First the positional embeddings for all nodes in the graph are computed. This involves running RWR for $\alpha$ anchor nodes for $\beta$ iterations (Eq.  6). This takes $\mathcal{O}(\alpha\times\beta)$ time. (2) Next, GNN is called to compute embeddings of node. In each layer of GNN a node $v\in\mathcal{V}$ aggregates message from $d$ neighbors where $d$ is the average degree of a node. This takes $\mathcal{O}(\mathcal{V})$ time. This operation is repeat for $L$ layers. Since $L$ is typically 1 or 2 hence we ignore this factor. (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq. 10. This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time. (4) Finally for the selected node, its partition has to be determined using eq. 12 and  13. This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score. Steps 2-4 are repeated for $T^{\prime}$ iterations. Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Since $|\mathcal{V}|\times d\approx|\mathcal{E}|$, hence complexity is $o(|\mathcal{E}|\times|\mathcal{V}|\big{)}$ Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.

### v3

@@ 4.3. Results on Inductivity to Partition Count
Figure 2. Results on the initial warm start and the final cut values obtained by NeuroCUT at $k{=}10$. It shows that our neural model NeuroCUT (Final) performs more accurate node and partition selection to optimize the objective function. Subsequently, there is a significant difference between the initial and final cut values.
As detailed in Section 3.2.2, the decoupling of parameter size and the number of partitions allows NeuroCUT to generalize effectively to an unseen number of partitions. In this section, we empirically analyze the generalization performance of NeuroCUT to an unknown number of partitions. We compare it with the non-neural baseline K-means, as the neural methods like GAP, DMon, Ortho, and MinCutPool cannot be employed to infer on an unseen number of partitions. In Table 7, we present the results on inductivity. Specifically, we trained NeuroCUT on different partition sizes ($k=5$ and $k=8$) and tested it on an unseen partition size $k=10$. The inductive version, referred to as NeuroCUT-I [^3] For clarity purposes, in this experiment we use NeuroCUT-T to refer to the transductive setting whose results are presented in Sec.  4.2, obtains high-quality results on different datasets. Note that while the non-neural method such as K-means needs to be re-run for the unseen partitions, NeuroCUT only needs to perform forward pass to produce the results on an unseen partition size.
4.4. Ablation Studies
Figure 3. Node Selection in Phase 1 with Heuristic vs Random: Relative % improvement (gain) in cut values obtained by NeuroCUT when using different node selection strategies for $k=5$. In most cases, our heuristic finds significantly better cuts than a random node selection procedure.
@@ 4.4. Ablation Studies
Impact of Clustering Initialization in Warm-start Phase: To understand the importance of different initializations in the warm-up phase(§ 3.1), in Appendix  A.6 we perform an ablation study using three different initialization schemes, namely K-means(default), density-based clustering DBSCAN (Ester et al., 1996), and Random initialization.
4.5. Impact of Cluster Strength on Performance
In Section A.8 in appendix, we analyze the impact of cluster strength in networks on the performance of different methods.
5. Conclusion
In this work, we study the problem of graph partitioning with node features. Existing neural methods for addressing this problem require the target objective to be differentiable and necessitate prior knowledge of the number of partitions. In this paper, we introduced NeuroCUT, a framework to effectively address the graph partitioning problem with node features. NeuroCUT tackles these challenges using a reinforcement learning-based approach that can adapt to any target objective function. Further, attributed to its decoupled parameter space and partition count, NeuroCUT can generalize to an unseen number of partitions. The efficacy of our approach is empirically validated through an extensive evaluation on four datasets, four graph partitioning objectives and diverse partition counts. Notably, our method shows significant performance gains when compared to the state-of-the-art techniques, proving its competence in both inductive and transductive settings.
Limitations and Future Directions: Achieving a sub-quadratic computational complexity with an inductive neural method for attributed graph partitioning is an open challenge. In NeuroCUT, node selection and perturbation are performed in a sequential fashion. One direction to improve the efficiency of NeuroCUT could be batch processing the selection and perturbation of multiple independent nodes simultaneously. Another interesting direction could be designing explanations of the neural methods—such as NeuroCUT—for graph combinatorial problems (Kakkad et al., 2023).
6. Acknowledgement
We thank Kartik Sharma for the helpful suggestions. Rishi Shah acknowledges that the funding for the conference was provided by the CMU GSA Conference Funding. Sahil Manchanda acknowledges financial support by GP Goyal grant of IIT Delhi and Qualcomm Innovation Fellowship India.
@@ Appendix A Appendix
A.1. Clustering for Initialization
@@
As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions. Towards this, we apply K-means algorithm (MacQueen et al., 1967) on the nodes of the graph where a node is represented by its raw node features and positional representation i.e $\mathbf{emb_{init}}(u)\;\forall u\in\mathcal{V}$ based upon eq. 8. Further, we used $L_{\inf}$ norm as the distance metric for clustering.
A.2. Time Complexity Analysis of NeuroCUT
@@
(1) First the positional embeddings for all nodes in the graph are computed. This involves running RWR for $\alpha$ anchor nodes for $\beta$ iterations (Eq.  6). This takes $\mathcal{O}(\alpha\times\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times|\mathcal{E}|})$ time. (2) Next, GNN is called to compute embeddings of node. In each layer of GNN a node $v\in\mathcal{V}$ aggregates message from $d$ neighbors where $d$ is the average degree of a node. This takes $\mathcal{O}(\mathcal{V})$ time. This operation is repeat for $L$ layers. Since $L$ is typically 1 or 2 hence we ignore this factor. (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq. 10. This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time. (4) Finally for the selected node, its partition has to be determined using eq. 12 and  13. This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score. Steps 2-4 are repeated for $T^{\prime}$ iterations. Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times|\mathcal{E}|})+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Since $|\mathcal{V}|\times d\approx|\mathcal{E}|$, hence complexity is $o(|\mathcal{E}|\times|\mathcal{V}|\big{)}$ Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
A.3. Time Complexity Comparison
Table 8 presents the complexities of NeuroCUT and other prominent neural and non-neural baselines algorithms. In the table, $|\mathcal{V}|$, $|\mathcal{E}|$, $k$, $d$ are the number of nodes, edges, partitions, and average node degree respectively. While some neural algorithms exhibit faster complexity, they require separate training for each partition size ($k$). In contrast, NeuroCUT, once trained, can generalize to any value. Consequently, for practical workloads, NeuroCUT presents a more scalable option in terms of computation overhead and storage (one model versus separate models for each).
Table 8. Time complexity of different methods.
Method Complexity Comments
DMon ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(d^{2}{\times}|\mathcal{V}|+|\mathcal{E}|)}$ Per iteration complexity
MinCutPool ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(d^{2}{\times}|\mathcal{V}|+|\mathcal{E}|)}$ Per iteration complexity
Ortho ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(|\mathcal{V}|{\times}k^{2})}$ Per iteration complexity
GAP ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(|\mathcal{V}|{\times}k^{2}+|\mathcal{V}|^{2})}$ Per iteration complexity
Spectral ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{O}(|\mathcal{V}|^{3})}$ Inference complexity
NeuroCUT ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}o(|\mathcal{V}|^{2})}$ Inference complexity
The quadratic time complexity of NeuroCUT may pose challenges for very large graphs. However, this complexity remains faster than spectral clustering, a widely used graph partitioning algorithm. Also, for the baseline neural methods(DMon, MinCutPool, Ortho and GAP), the time complexity provided is per iteration. The number of iterations often ranges between 1000 to 2000 and there is no clear understanding of how it varies as a function of the graph (like density, diameter, etc.)
A.4. Comparison: Non-neural Baselines
We compare the performance of our method against graph based clustering algorithms hMETIS and Spectral Clustering and partition algorithm Gomory-Hu Tree (Gomory and Hu, 1961). Since these methods are not compatible with datasets having raw node features, we compare with two datasets namely Facebook (Leskovec and Mcauley, 2012) and Stochastic Block Model(SBM) (Abbe, 2017) which don’t have node features for this comparision. Table 9 compares the performance of NeuroCUT against hMETIS and Spectral Clustering. The details of these datasets are presented in Table 2. In addition to non-neural methods, we also show the performance of neural methods MinCutPool, DMon, Ortho and GAP on these datasets.
The performance of NeuroCUT is better than non-neural methods hMETIS, Spectral and Gomory-Hu Tree in most of the cases on Facebook dataset. Further, on SBM dataset, it matches the performance of hMETIS and Spectral clustering. The SBM dataset has a clear community structure, making it an easy instance for partitioning algorithms, resulting in similar performance across different methods. We would also like to highlight that in the case of $k$--$mincut$, Gomory-Hu Tree algorithm generated trivial partitions for these datasets where almost all nodes were assigned the same partition. The other neural methods such as MinCutPool, DMon, Ortho, and GAP fail to produce a valid solution in most of the cases. This is possibly because these baselines are not robust to perform on datasets without raw node features. Overall, NeuroCUT outperforms the baselines on diverse objectives and datasets.
Table 9. Performance of different methods on Facebook and SBM dataset. Lower values are better. ${}^{\prime}-^{\prime}$ denotes nan.
Dataset Method Metrics
Normalized Sparsest Balanced $k$-MinCut
Facebook NeuroCUT 0.257 4.121 0.972 0.015
hMETIS 1.15 52.319 1.115 0.200
Spectral 1.67 4.72 2.459 0.003
Gomory-Hu Tree 4.00 5.00 4.792 $\sim$0
MinCutPool - - - 0.023
DMon - - - $\sim$0
Ortho - - - 0.05
GAP - - - 0.025
SBM NeuroCUT 0.191 3.939 0.191 0.038
hMETIS 0.191 3.939 0.191 0.038
Spectral 0.191 3.939 0.191 0.038
Gomory-Hu Tree 4.003 48.75 4.787 0.0075
MinCutPool - - - 0.0
DMon - - - 0.0
Ortho - - - $\sim$0
GAP - - - 0.035
A.5. Comparison with Gatti et al. (Gatti et al., 2022)
In Table 10, we compare the performance of our method against DRL method proposed by Gatti et al. (Gatti et al., 2022) which solves the normalized cut problem for the case where the number of partitions is exactly two. We observe that NeuroCUT outperforms DRL on all datasets.
Table 10. Gatti et al. on Normalized Cut at $k=2$.
$\dfrac{\text{\bf Method}\rightarrow}{\text{\bf Dataset}\downarrow}$ Cora CiteSeer Harbin Actor SBM Facebook
NeuroCUT 0.02 0.02 0.01 0.17 0.191 0.257
Gatti et al. 0.355 0.29 0.13 1.00 0.72 0.92
A.6. Impact of Clustering Initialization
To understand the importance of different initializations in the warm-up phase, we perform an ablation study on $2$ datasets namely Cora and CiteSeer using 3 different initialization namely K-means, density-based clustering DBSCAN (Ester et al., 1996), and Random initialization. In Table 11 we observe the performance on normalized cut at $k=5$. We observe that K-means performs the best in this experiment. The improvement observed when using K-means or DBSCAN over Random shows that a good initialization i.e. warm-up does help in improving quality of partitions. In DBSCAN we set $eps=0.9$ and $min\_samples=2$. It is worth noting that DBSCAN’s performance can vary based on the parameter selection. Nonetheless, the primary goal of this experiment is to demonstrate the advantageous role of a good initialization, specifically in contrast to random initialization.
Table 11. Impact of Clustering initialization in Warm-up Phase of NeuroCUT in the normalized cut objective at $k=5$.
Dataset K-means Random DBSCAN
Cora 0.33 0.79 0.55
CiteSeer 0.20 0.51 0.70
Table 12. Statistics for Normalized Cut (3 runs) for $k=5$
Dataset Our Heuristic Random Selection Random Selection % Gain
Avg Best (from best)
Cora 0.33 $\pm$ 0.0024 0.461 $\pm$ 0.0164 0.445 35
Citeseer 0.20 $\pm$ 0.00961 0.276 $\pm$ 0.0177 0.258 29
Harbin 0.073 $\pm$ 0.0027 0.097 $\pm$ 0.0131 0.086 18
Table 13. Statistics for Balanced Cut (across 3 runs) for $k=5$
Dataset Our Heuristic Random Selection Random Selection % Gain
Avg Best (from best)
Cora 0.642 $\pm$ 0.0065 0.753 $\pm$ 0.0574 0.672 2
Citeseer 0.248 $\pm$ 0.0049 0.307 $\pm$ 0.039 0.252 0.5
Harbin 0.1087 $\pm$ 0.0010 0.1313 $\pm$ 0.011 0.1293 19
Table 14. Number of Nodes Re-assigned after Initial Assignment by NeuroCUT on the Balanced Cut objective.
Dataset Nodes Perturbed Total Nodes Percentage
Cora 791 2495 31.7
Citeseer 413 2120 19.4
Harbin 187 6598 2.8
A.7. Impact of Node Selection Procedures
In Fig. 3 in main paper, random selection performs well only in Balanced Cuts. The objective in Balanced Cuts (Eq. 3.) is not only a function of the combined weight of cut edges, but also balancing the number of nodes across partitions. This additional node-balancing term is not present in the objective functions of the other cut definitions. Due to this reduced importance of optimizing the cut value, random does well, since even when an incorrect node is selected, it can still be utilized to keep the partition sizes balanced. Moreover, we find that when the warm-up assignment is less accurate and the percentage of nodes re-assigned by NeuroCUT to a different partition is high, a random selection of nodes is closer in efficacy to our heuristic selection. In contrast, in Harbin, the percentage of nodes perturbed to a new partition is significantly smaller. In this scenario, the probability that random selection will select this small subset is lower and hence amplifying the effectiveness gap between random and heuristic selection.
A.8. Impact of Cluster Strength on Performance
To understand the impact of community structure on performance of NeuroCUT, we generate Stochastic Model Block(SBM) graphs with different intra cluster strength. In Table 15, we observe the performance of different methods on the normalized cut objective at $k=5$. The baseline methods hMETIS and Spectral Clustering perform worse when the strength within communities is low indicated by Intra Cluster Edge Probability and Clustering Coefficient values. Although NeuroCUT outperforms or matches existing methods in all cases, however, the gap between baselines and NeuroCUT increases significantly when the community structure is not strong. Further, the neural baselines DMon, MinCutPool and Ortho generated ‘nan’ values for this experiment.
Table 15. Performance on SBM Dataset
SBM Dataset Statistics Normalized Cut
Intra Cluster Edge Probability Clustering Coefficient $|\mathcal{V}|$ $|\mathcal{E}|$ NeuroCUT hMETIS Spectral Clustering
1 0.005 0.003 495 555 0.3089 0.331 0.379
2 0.01 0.0029 525 661 0.3434 0.5889 0.4259
3 0.2 0.184 500 5150 0.1912 0.1912 0.1912
4 0.4 0.3822 500 10102 0.1153 0.1153 0.1153
A.9. Impact of $\beta$ on NeuroCUT
We study the impact of $\beta$ parameter(eq. 6) which is the number of iterations in random walk with restart. In Table 16 we present the performance of NeuroCUT using different $\beta$ on normalized-cut objective at $k=5$. We observe that NeuroCUT improves with more iterations as $\beta$ increases and then its performance stabilizes.
Table 16. Impact of $\beta$ parameter on NeuroCUT on normalized cut at $k=5$ on the Cora dataset.
$\beta$ Normalized Cut
1 2.12
3 1.57
10 0.33
50 0.33
100 0.33
A.10. Stability Across Multiple Runs
Table 17 shows the Mean and Standard deviation of NeuroCUT for Normalized cut at $k=5$ for 3 runs. The Std Dev. is lower than the Mean showing the stability of NeuroCUT across multiple runs.
Table 17. Mean and Standard Deviation for different datasets on Normalized Cut at $k=5$ for NeuroCUT
Dataset Mean $\pm$ Std Dev
Cora $0.337\pm 0.0024$
Citeseer $0.20\pm 0.00961$
Harbin $0.073\pm 0.0027$
Actor $0.97\pm 0.0389$